OpenStudy (anonymous):

A landscape architect wishes to enclose a rectangular garden on one side by a brick wall costing $21 per foot and on the other three sides by a metal fence costing $7 per foot. If the area of the garden is 800 ft, find the dimensions of the garden minimizing the cost. (Let x be the length of the brick wall and y be the length of an adjacent side in feet.)

OpenStudy (anonymous):

here are two equations that can be figured out from the question 21*x + 7*y = total cost x*y = 800 can you figure out how to relate them into one equation, in order to solve for the minimum cost?

OpenStudy (gorv):



OpenStudy (gorv):

length of the brick wall =x length of an adjacent side =y area=x*y=800 x*y=800 brick wall cost=$21*x metal fence cost=$7*y+$7y+*7x total cost=$28*x+$14*y x*y=800 y=800/x total cost=$28*x+$14*800/x total cost=28x+11200/x differntiate d(total cost)/dx =28-11200/x^2 =(28*x^2-11200)/x^2 tofind value of x put equation =0 (28*x^2-11200)/x^2=0 (28*x^2-11200)=0 x^2=11200/28=400 x=20 now find d^2(total cost)/dx^2 =-11200* (-2)/x^3 now calculate its value at x=20 =22400/20^3=22400/8000 >0 if second derivative is > 0 then function will have minimum value at x=20 x=20feet x*y=800 y=800/20=40feet x= 20 feet y=40 feet

OpenStudy (gorv):

@marjorie